OpenStudy (adamsaleh):

(1- sinxtanx)/1+secx=cosx

OpenStudy (vishweshshrimali5):

@adamsaleh , is this your question: \(\large{\cfrac{1-\sin{x}\tan{x}}{1+\sec{x}} = \cos{x}}\) Do we have to prove it or solve for solutions of x ?

OpenStudy (gudden):

@adamsaleh .. you there???

OpenStudy (gudden):

Do you know that we can write tan in terms of cos and sin? If yes, then show that ....

OpenStudy (anonymous):

\[1-\sin ^{2x ^{}}/cosx\]=cosx+1

OpenStudy (anonymous):

1\[cosx-\sin ^{2 ^{}}x=\cos ^{2}x ^{}+1\]